[Brahms-l] Shift report 20040401 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Thu Apr 01 2004 - 07:45:02 EST
------------------------------------------------------------
Date: Thu Apr  1 07:45:01 2004, Shift: 00:00-08:00
Supervisor: sl, with: 
Log book#15 pages 104 - 104

Shift Summary:
------------------------------------------------------------
1.No beam until late morning, Now there is a problem with the yellow beam that looses intensity very fast.
2.Only pedestal runs excpet one short data run 11532 with bad beam.
3.T2 B anode tripped.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Apr 1 07:45:28 2004

This archive was generated by hypermail 2.1.8 : Thu Apr 01 2004 - 07:45:51 EST